# Artificial barriers for stochastic differential equations and for construction of boundary-preserving schemes

### v2

@@ 4.1. Construction of artificial barriers
The second step of the method of artificial barriers is to augment the SDE (19) with what we call artificial barriers. To this end, we consider the following process $\xi_{N}$ with reflection (see Section 3.2 for details) that can be viewed as $X_{N}$ augmented with artificial barriers:
[…]
@@
with initial value $U_{N}(0)=0$, is the upper barrier at $b-1/N$ and where
@@
$\xi_{N}(t)=x_{0}+\int_{0}^{t}f_{N}(\xi_{N}(s))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g_{N}(\xi_{N}(s))\mathop{}\!\mathrm{d}B(s)-U_{N}(t)+L_{N}(t)$
@@
(24) $w_{N}(t)=x_{0}+\int_{0}^{t}f_{N}(\xi_{N}(s))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g_{N}(\xi_{N}(s))\mathop{}\!\mathrm{d}B(s),$
@@
(25) ${\mathbb{E}}\left[\sup_{t\in[0,T]}|\xi_{N}(t)|^{p}\right]\leq\begin{cases}C(p,T,L_{f},L_{g},a,x_{0}),\text{ if }b=\infty,\\ \max(|a|^{p},|b|^{p}),\text{ if }|b|<\infty,\end{cases}$
@@
Moreover, the constant $C(p,T,L_{f},L_{g},a,b,x_{0})$ is independent of $N\in\mathbb{N}$. If $b=\infty$, then the constant for the time increments is independent of $b$.
@@
The following proposition is the reason why we call the addition of the processes $U_{N}$ and $L_{N}$ artificial barriers, as they do not change the process $X_{N}$.
@@
Suppose Assumptions 1 and 2 are satisfied and let $N\in\mathbb{N}$. Then the processes $X_{N}$ solving (19) and $\xi_{N}$ solving (23) are equal almost surely:
@@
If $X_{N}$ is a solution of (19) and $X_{N}\in D_{N}$, then $(X_{N},\ U_{N}\equiv 0,L_{N}\equiv 0)$ is a solution to (23). By uniqueness of solutions of (23), we conclude that $X_{N}$ and $\xi_{N}$ are equal almost surely. ∎
The third step of the method of artificial barriers is to discretise the RSDE (23) using any (non-strict) boundary-preserving numerical scheme for reflected SDEs. To this end, we introduce a discretisation parameter $M\in\mathbb{N}$ and we discretise the time interval $[0,T]$ into $M\in\mathbb{N}$ intervals $[t_{m},t_{m+1}]$ each of size $\Delta t=T/M$. We also let $\ell_{\Delta t}(s)=t_{k}$ for $s\in[t_{k},t_{k+1})$ that will be used in Section 5 and in Section 6. Then, in particular, $t_{m}=m\Delta t$ for every $m=0,\ldots,M$. For now, to complete the convergence statement in this section, we assume that $\hat{\xi}_{N,\Delta t}$ is a (non-strict) boundary-preserving scheme for the RSDE in (23) with error constant independent of $N\in\mathbb{N}$.
@@
Let $N\in\mathbb{N}$ and $\Delta t>0$. For any $x_{0}\in D_{1}$, we can find an approximation $\hat{\xi}_{N,\Delta t}$ to the solution $\xi_{N}$ of the RSDE in (23) that only takes values in $\operatorname{cl}D_{N}$
@@
In the next sections we define the Artificial Barriers EM (ABEM) scheme (Section 5) and the Artificial Barriers Euler–Peano (ABEP) scheme (Section 6) and we verify that both satisfy Assumption 3. The ABEM and ABEP schemes combine artificial barriers with a modified version of the classical projected EM scheme for RSDEs and with a modified version of the classical Euler–Peano scheme for RSDEs, respectively. Note that the method of artificial barriers can be combined with other boundary-preserving schemes for RSDEs (for example, the projected and reflected scheme, see [26]). We now state the main result of this section.
@@
for every $p\geq 2$, where the constant is independent of $N$ and $\Delta t$. If $b=\infty$, then the constant for the $L^{p}(\Omega)$-error is independent of $b$.
@@
Let $p\geq 2$. The first statement follows from $\operatorname{cl}D_{N}\subset D$, for every $N\in\mathbb{N}$, and from Assumption 3. Recall the definitions of $X_{N}$ and $\xi_{N}$ in equation (19) and in equation (23), respectively. Then, by the triangle inequality, we have that
@@
$\sup_{t\in[0,T]}|\hat{\xi}_{N,\Delta t}(t)-X(t)|\leq\eta_{\epsilon}\Delta t^{\gamma-\epsilon}$
almost surely. Moreover, $\eta_{\epsilon}$ is independent of $N\in\mathbb{N}$.
@@ 5. Artificial Barriers Euler–Maruyama
First, we combine the method of artificial barriers from Section 4 with a modification of the classical projected Euler–Maruyama (PEM) scheme for RSDEs applied to the RSDE with modified coefficients in (23). We call the resulting scheme the Artificial Barriers EM scheme and abbreviate it as ABEM. We first introduce the ABEM scheme and then we show that the ABEM scheme satisfy Assumption 3 in Section 4. It will then follow that the scheme is boundary-preserving and converges to the solution $X$ of the considered SDE (1), by Theorem 22. We remark that the standard projected EM (PEM) scheme applied to the SDE in (1) is non-strict boundary-preserving, meaning that the approximations might hit the boundary $\partial D$. Non-strict boundary-preservation is particularly problematic in our considered cases where both $f$ and $g$ typically have zeros at the boundary, since such schemes would forever remain at the boundary once it hits it.
$L^{p}(\Omega)$-convergence of order $1/2-$ of the PEM scheme for RSDEs is a known result in the literature (see, for example, [33] for the case $p=2$). We show in this section that this is also true for the ABEM scheme for the RSDEs in (23) and that the $L^{p}(\Omega)$-error constant can be taken to be independent of $N$, and hence the ABEM scheme satisfy Assumption 3.
Let $\Pi_{N}:\mathbb{R}\to\operatorname{cl}D_{N}$ denote the projection onto $\operatorname{cl}D_{N}$ defined by
@@
Recall that $x_{0}\in D_{1}$. We define the ABEM scheme, denoted by $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}$, as follows: Let $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(0)=x_{0}$ and successively define $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k+1})$ as
(28) $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k+1})=\Pi_{N}\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k})+f\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k})\right)\Delta t+g\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k})\right)\Delta B_{k}\right),$
where $\Delta B_{k}=B(t_{k})-B(t_{k-1})$. Note that if we would use the modified coefficient functions $f_{N}$ and $g_{N}$ instead of $f$ and $g$ in (28), then the scheme could get stuck at the boundary points of $\operatorname{cl}D_{N}$. Also note that equation (28) can equivalently be written as
(29) $\begin{split}\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k+1})=\min\{b-1/N,\max\{&a+1/N,\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k})\\ &+f(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k}))\Delta t+g(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k}))\Delta B_{k}\}\}.\end{split}$
@@
For the purpose of the convergence analysis, we extend the definition in equation (28) to continuous time and we introduce an auxiliary process $\hat{w}_{N,\Delta t}^{\textrm{ABEM}}$: For $t\in[t_{k},t_{k+1})$, let
(30) $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t)=\Pi_{N}\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k-1})+f\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k-1})\right)\Delta t+g\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k-1})\right)\Delta B_{k}\right),$
@@
$\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k-1})+f\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k-1})\right)\Delta t+g\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{k-1})\right)\Delta B_{k}.$
@@
(31) $\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t)=x_{0}+\int_{0}^{t}f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}B(s)$
that coincides with $\hat{w}_{N,\Delta t}^{\textrm{ABEM}}$ on the grid points $\{t_{k}:\ k=0,\ldots,M\}$: $\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k})=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k})$ for $k=0,\ldots,M$. Indeed by an induction argument over the index $k$: Suppose that $\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k})=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k})$ for some $k=0,\ldots,M-1$. Then
$\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k+1})$ $=\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k})+\int_{t_{k}}^{t_{k+1}}f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}s+\int_{t_{k}}^{t_{k+1}}g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}B(s)$
$=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k})+f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{k}))\Delta t+g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{k}))\Delta B_{k+1}$
$=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k+1}),$
where we used that $\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}$ is constant on $[t_{k},t_{k+1})$. Moreover,
@@
with $\tilde{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t)=\Gamma_{N}(\tilde{w}_{N,\Delta t}^{\textrm{ABEM}})(t)$ for $t\in[0,T]$, is an associated triple on $\operatorname{cl}D_{N}$ in the sense of Definition 8.
@@
${\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t)\right|^{p}\right]={\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\Gamma_{N}\left(\hat{w}_{N,\Delta t}^{\textrm{ABEM}}\right)(t)\right|^{p}\right]\leq 2^{p}{\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)\right|^{p}\right]+|a|^{p}+1$
@@
$\sup_{t\in[0,r]}\left|\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)\right|^{p}=\sup_{k=0,\ldots,\ell_{\Delta t}(r)/\Delta t}\left|\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k})\right|^{p}$ $=\sup_{k=0,\ldots,\ell_{\Delta t}(r)/\Delta t}\left|\tilde{w}^{\textrm{ABEM}}_{N,\Delta t}(t_{k})\right|^{p}$
@@
where we recall that $\ell_{\Delta t}(s)=t_{k}$ for $s\in[t_{k},t_{k+1})$ and where we also used that $\hat{w}_{N,\Delta t}^{\textrm{ABEM}}$ and $\tilde{w}^{\textrm{ABEM}}_{N,\Delta t}$ agree on the time grid points (see paragraph before this proposition). By (31), we can bound
@@
For the Itô integral, we use the BDG inequality and linear growth of $g$ on $D$ (see Assumption 1) and of $\Gamma_{N}$ (see (15)) to estimate
@@
We verify that Assumption 3 is fulfilled and the result then follows from Theorem 22. The proof follows the same proof idea as in [33] with the extra issue that the coefficient functions of the reflected SDE for $\xi_{N}$ in (23) are not the same as in the definition of the Artificial Barriers EM scheme $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}$. In view of Proposition 24, it remains to prove that

### v3

@@ 4.1. Construction of artificial barriers
The second step of the method of artificial barriers is to augment the SDE (19) with what we call artificial barriers. To this end, we consider the following reflected process $\xi_{N}$ (see Section 3.2 for details) that can be viewed as $X_{N}$ augmented with artificial barriers:
(24) $\left\{\begin{aligned} &\mathop{}\!\mathrm{d}\xi_{N}(t)=f_{N}(\xi_{N}(t))\mathop{}\!\mathrm{d}t+g_{N}(\xi_{N}(t))\mathop{}\!\mathrm{d}B(t)-\mathop{}\!\mathrm{d}U_{N}(t)+\mathop{}\!\mathrm{d}L_{N}(t),\ t\in(0,T],\\ &\xi_{N}(0)=x_{0}\in D_{N},\end{aligned}\right$.
@@
with initial value $U_{N}(0)=0$, is the upper barrier at b1/Nb-1/N and where
@@
$\xi_{N}(t)=x_{0}+\int_{0}^{t}f_{N}(\xi_{N}(s))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g_{N}(\xi_{N}(s))\mathop{}\!\mathrm{d}B(s)-U_{N}(t)+L_{N}(t),\ t\in[0,T],$
@@
(25) $w_{N}(t)=x_{0}+\int_{0}^{t}f_{N}(\xi_{N}(s))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g_{N}(\xi_{N}(s))\mathop{}\!\mathrm{d}B(s),\ t\in[0,T],$
@@
(26) ${\mathbb{E}}\left[\sup_{t\in[0,T]}|\xi_{N}(t)|^{p}\right]\leq\begin{cases}C(p,T,L_{f},L_{g},x_{0}),\text{ if }b=\infty,\\ \max(|a|^{p},|b|^{p}),\text{ if }|b|<\infty,\end{cases}$
@@
Moreover, the constants are independent of $N\in\mathbb{N}$. If $b=\infty$, then the constant for the time increments is independent of $a$ and $b$.
@@
The following proposition is the reason we call the processes $U_{N}$ and $L_{N}$ artificial barriers, as their inclusion have no impact on the process $X_{N}$.
@@
Suppose Assumptions 1 and 2 are satisfied and let $N\in\mathbb{N}$. Then the solution $X_{N}$ of (19) and the solution $\xi_{N}$ of (24) are, almost surely, equal:
@@
If $X_{N}$ is a solution of (19) and $X_{N}\in D_{N}$, then $(X_{N},\ U_{N}\equiv 0,L_{N}\equiv 0)$ is a solution to (24). By uniqueness of solutions of (24), we conclude that $X_{N}$ and $\xi_{N}$ are equal, almost surely. ∎
The third step of the method of artificial barriers is to discretise the RSDE (24) using any (non-strict) boundary-preserving numerical scheme for RSDEs. To this end, we introduce a discretisation parameter $M\in\mathbb{N}$ and we discretise the time interval $[0,T]$ into $M\in\mathbb{N}$ intervals $t_{m},t_{m+1}]$ each of size $\Delta t=T/M$. We also let $\ell_{\Delta t}(s)=t_{m}$ for $s\in[t_{m},t_{m+1})$ that will be used in Section 5 and in Section 6. Then, in particular, $t_{m}=m\Delta t$ for every $m=0,\ldots,M$. For now, to complete the convergence statement in this section, we assume that $\hat{\xi}_{N,\Delta t}$ is a (non-strict) boundary-preserving scheme for the RSDE in (24) with error constant independent of $N\in\mathbb{N}$.
@@
Let $N\in\mathbb{N}$ and $\Delta t>0$. For every $x_{0}\in D_{1}$, there exists an approximation $\hat{\xi}_{N,\Delta t}$ to the solution $\xi_{N}$ of the RSDE in (24) that only takes values in $\operatorname{cl}D_{N}$
@@
In the next sections we define the Artificial Barriers EM (ABEM) scheme (Section 5) and the Artificial Barriers Euler–Peano (ABEP) scheme (Section 6), and we verify that both satisfy Assumption 3. The ABEM and ABEP schemes combine artificial barriers with a modified version of the classical projected EM scheme for RSDEs and with a modified version of the classical Euler–Peano scheme for RSDEs, respectively. Note that the method of artificial barriers can be combined with other boundary-preserving schemes for RSDEs (for example, the projected and reflected scheme, see [33]). We now state the main result of this section.
@@
for every $p\geq 2$, where the constant is independent of $N$ and $\Delta t$. If $b=\infty$, then the constant is independent of $b$.
@@
Let $p\geq 2$. The first statement follows from $\operatorname{cl}D_{N}\subsetneq D$, for every $N\in\mathbb{N}$, and from Assumption 3. Recall the definitions of $X_{N}$ and $\xi_{N}$ in equation (19) and in equation (24), respectively. Then, by the triangle inequality, we have that
@@
$\sup_{t\in[0,T]}|\hat{\xi}_{N,\Delta t}(t)-X(t)|\leq\eta_{\epsilon}\Delta t^{\gamma-\epsilon},$
almost surely. Moreover, $\eta_{\epsilon}$ is independent of $N\in\mathbb{N}$ and $\Delta t$.
@@ 5. Artificial Barriers Euler–Maruyama
In this section, we combine the method of artificial barriers from Section 4 with a modification of the classical projected Euler–Maruyama (PEM) scheme for RSDEs applied to the RSDE with modified coefficients in (24). We call the resulting scheme the Artificial Barriers EM scheme and abbreviate it as ABEM. We first introduce the ABEM scheme, and then we show that the ABEM scheme satisfy Assumption 3 in Section 4. We can from this conclude that the scheme is boundary-preserving and converges to the solution $X$ of the considered SDE (1), by Theorem 22. We remark that the standard projected EM (PEM) scheme applied to the SDE in (1) is non-strict boundary-preserving, meaning that the approximations might hit the boundary $\partial D$. A non-strict boundary-preserving numerical scheme is particularly problematic in our considered cases where both $f$ and $g$ typically have zeros at the boundary, since such a scheme would forever remain at the boundary once it hits it.
$L^{p}(\Omega)$-convergence of order $1/2$- of the PEM scheme for RSDEs is a known result in the literature (see, for example, [41] for the case $p=2$). We show in this section that this is also true for the ABEM scheme for the RSDEs in (24) and that the constants for the moment bounds and the $L^{p}(\Omega)$-error can be taken to be independent of $N$ and $\Delta t$, and hence the ABEM scheme satisfy Assumption 3.
Let $\Pi_{N}:\mathbb{R}\to\operatorname{cl}D_{N}$ denote the projection onto $\operatorname{cl}D_{N}$, defined by
@@
If $b=\infty$, then we interpret x>b1/Nx>b-1/N in (28) as never true. We define the ABEM scheme, denoted by $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}$, as follows: Let $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(0)=x_{0}\in D_{1}$ and successively define $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m+1})$, for $m=0,\ldots,M-1$, as
(29) $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m+1})=\Pi_{N}\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m})+f\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m})\right)\Delta t+g\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m})\right)\Delta B_{m}\right),$
where $\Delta B_{m}=B(t_{m+1})-B(t_{m})$. Replacing $f$ and $g$ in (29) with $f_{N}$ and $g_{N}$, respectively, would make it possible for scheme to get stuck at the boundary points of $\operatorname{cl}D_{N}$. Also note that equation (29) can equivalently be written as
(30) $\begin{split}\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m+1})=\min\{b-1/N,\max\{&a+1/N,\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m})\\ &+f(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m}))\Delta t+g(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m}))\Delta B_{m}\}\}.\end{split}$
@@
For the purpose of the convergence analysis, we extend the definition in equation (29) to continuous time, and we introduce an auxiliary process $\hat{w}_{N,\Delta t}^{\textrm{ABEM}}$: For $t\in[t_{m},t_{m+1})$, let
(31) $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t)=\Pi_{N}\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m-1})+f\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m-1})\right)\Delta t+g\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m-1})\right)\Delta B_{m-1}\right),$
@@
$\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m-1})+f\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m-1})\right)\Delta t+g\left(\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t_{m-1})\right)\Delta B_{m-1}.$
@@
(32) $\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t)=x_{0}+\int_{0}^{t}f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}B(s),\ t\in[0,T],$
that coincides with $\hat{w}_{N,\Delta t}^{\textrm{ABEM}}$ on the grid points $\{t_{m}:\ m=0,\ldots,M\}$: $\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m})=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m})$ for $m=0,\ldots,M$. Indeed by an induction argument over the index $m$: Suppose that $\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m})=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m})$ for some $m=0,\ldots,M-1$. Then
$\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m+1})$ $=\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m})+\int_{t_{m}}^{t_{m+1}}f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}s+\int_{t_{m}}^{t_{m+1}}g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}B(s)$
$=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m})+f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{m}))\Delta t+g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{m}))\Delta B_{m}$
$=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m+1}),$
where we used that $\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}$ is constant on $t_{m},t_{m+1})$. Moreover,
@@
with $\tilde{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t)=\Gamma_{N}(\tilde{w}_{N,\Delta t}^{\textrm{ABEM}})(t)$, for $t\in[0,T]$, is an associated triple on $\operatorname{cl}D_{N}$ in the sense of Definition 8.
@@
${\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t)\right|^{p}\right]={\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\Gamma_{N}\left(\hat{w}_{N,\Delta t}^{\textrm{ABEM}}\right)(t)\right|^{p}\right]\leq C(p)\left({\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)\right|^{p}\right]+|a|^{p}+1\right)$
@@
$\sup_{t\in[0,r]}\left|\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)\right|^{p}=\sup_{m=0,\ldots,\ell_{\Delta t}(r)/\Delta t}\left|\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m})\right|^{p}$ $=\sup_{m=0,\ldots,\ell_{\Delta t}(r)/\Delta t}\left|\tilde{w}^{\textrm{ABEM}}_{N,\Delta t}(t_{m})\right|^{p}$
@@
where we recall that $\ell_{\Delta t}(s)=t_{m}$ for $s\in[t_{m},t_{m+1})$ and where we also used that $\hat{w}_{N,\Delta t}^{\textrm{ABEM}}$ and $\tilde{w}^{\textrm{ABEM}}_{N,\Delta t}$ agree on the time grid points (see paragraph before this proposition). By (32), we can bound
@@
For the Itô integral, we use the BDG inequality, Jensen’s inequality for integrals, and linear growth of $g$ on $D$ (see Assumption 1) and of $\Gamma_{N}$ (see (15)) to estimate
@@
We verify that Assumption 3 is fulfilled and the result then follows from Theorem 22. The proof follows the same proof idea as in [41] with the extra issue that the coefficient functions of the RSDE for $\xi_{N}$ in (24) are not the same as in the definition of the Artificial Barriers EM scheme $\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}$. In view of Proposition 24, it remains to prove that